# Risk-Averse Learning with Delayed feedback

### v1

@@ 3.1 One-point risk-averse learning
with probability $1-\gamma$. When $a<1$, selecting $\delta=T^{-\frac{a}{4}}$ and $\eta=(T+D)^{-\frac{3a}{4}}$, Algorithm 1 achieves ${\rm R}_{1}(T)=\tilde{\mathcal{O}}\Big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{3a}{4}}+(T+D)^{-\frac{3a}{4}}\big{(}T^{1+\frac{a}{2}}+T^{\frac{a}{4}}D\big{)}\Big{)}$ with probability $1-\gamma$. Since $T^{1-\frac{a}{4}}\geq(T+D)^{-\frac{3a}{4}}T^{1+\frac{a}{2}}$, ${\rm R}_{1}(T)$ can be further simplified as in Theorem 1. When $a\geq 1$, selecting $\delta=T^{-\frac{1}{4}}$ and $\eta=(T+D)^{-\frac{3}{4}}$, Algorithm 1 achieves ${\rm R}_{1}(T)=\tilde{\mathcal{O}}\Big{(}(T+D)^{\frac{3}{4}}\Big{)}$ with probability $1-\gamma$. $\blacksquare$
Remark 1.
When selecting $\bar{\delta}=(T+D)^{-\frac{1}{4}\min\{1,a\}}$ and $\bar{\eta}=(T+D)^{-\frac{3}{4}\min\{1,a\}}$, Algorithm 1 achieves $\bar{{\rm R}}_{1}(T)=(T+D)^{1-\frac{1}{4}\min\{1,a\}}$ with probability $1-\gamma$, offering a concise but conservative regret bound for ${\rm R}_{1}(T)$. The explanation follows. When $a>1$, $\bar{{\rm R}}_{1}(T)={\rm R}_{1}(T)=\tilde{\mathcal{O}}\big{(}(T+D)^{\frac{3}{4}}\big{)}$. When $a<1$, for terms in (3.1), we have $\frac{D_{x}^{2}}{\eta}=\tilde{\mathcal{O}}\big{(}(T+D)^{\frac{3a}{4}})\big{)}$, $(3+D_{x}/r)\delta L_{0}T=\tilde{\mathcal{O}}\big{(}T^{1-\frac{a}{4}}\big{)}$, $\frac{cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2L_{T}/\gamma)}{2}}T^{1-\frac{a}{2}}=\tilde{\mathcal{O}}\big{(}T^{1-\frac{a}{4}}\big{)}$, and $\frac{\eta L_{0}dU}{\delta}D+\frac{\eta}{2}\frac{d^{2}U^{2}}{\delta^{2}}T\leq\frac{\eta L_{0}dU}{\delta}D+\frac{\eta}{2}\frac{d^{2}U^{2}}{\delta}T=\tilde{\mathcal{O}}\big{(}(T+D)^{1-\frac{3a}{4}}T^{\frac{a}{2}}\big{)}$. It can be observed that $(T+D)^{\frac{3a}{4}}$, $T^{1-\frac{a}{4}}$ and $(T+D)^{1-\frac{3a}{4}}T^{\frac{a}{2}}$ are all smaller than $(T+D)^{1-\frac{a}{4}}$ when $a<1$.
Theorem 1 analyzes the regret bound concerning the horizon length $T$ and cumulative delay $D$. The following proposition and corollary discuss the conditions on cumulative delay $D$ and maximum individual delay $\bar{d}$ required for the regret to be sublinear in $T$, respectively.
Proposition 1.
Let Assumptions 1, 2, 3 hold. When the cumulative delay satisfying $D\leq c_{1}T^{\alpha}$, where $c_{1}>0$, $\alpha<\beta$ and
${\beta}=\left\{\begin{array}[]{ll}\frac{4}{4-a}&{\rm if}~{}a\in(0,1)\\ \frac{4}{3}&{\rm if}~{}a\geq 1,\end{array}\right.$ (22)
Algorithm 1 achieves sublinear regret with probability $1-\gamma$.
Proof. We start from $a\geq 1$. Substituting $D=T^{\alpha}$ into the regret bound $\bar{{\rm R}}_{1}(T)=\tilde{\mathcal{O}}\Big{(}(T+D)^{\frac{3}{4}}\Big{)}$, ${\rm R}_{1}(T)$ is sublinear when $\frac{3\alpha}{4}<1$. Thus, we obtain the last line of (22). When $10<a<$ 1, we substitute $D=T^{\alpha}$ into $\bar{{\rm R}}_{1}(T)=\tilde{\mathcal{O}}\Big{(}(T+D)^{1-\frac{a}{4}}\Big{)}$. When $\alpha<1$, the sublinearity of $\bar{{\rm R}}_{1}(T)$ is straightforward. When $\alpha\geq 1$, for $\bar{{\rm R}}_{1}(T)$ to remain sublinear, it requires $(1-\frac{a}{4})\alpha<1$. This leads to the first line of (22). $\blacksquare$
Corollary 1.
Let Assumptions 1, 2, 3 hold. When the maximum delay satisfies $\bar{d}\leq c_{2}T^{\tilde{\alpha}}$, where $c_{2}>0$, $\tilde{\alpha}<\tilde{\beta}$ and
$\tilde{\beta}=\left\{\begin{array}[]{ll}\frac{a}{4-a}&{\rm if}~{}a\in(0,1)\\ \frac{1}{3}&{\rm if}~{}a\geq 1,\end{array}\right.$
Algorithm 1 achieves sublinear regret with probability $1-\gamma$.
Proof. To recap, $D=\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}+\bar{d}L_{T}-\frac{\bar{d}^{2}+\bar{d}}{2}$. Let $\bar{d}=T^{\tilde{\alpha}}$, we have $\bar{d}L_{T}\leq\bar{d}T=T^{1+\tilde{\alpha}}$ and $\bar{d}L_{T}\geq\bar{d}(T-\bar{d})=T^{1+\tilde{\alpha}}-T^{2\tilde{\alpha}}$. Additionally, we have $\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}\leq\bar{d}(L_{T}-2\bar{d}-1)\leq T^{1+\tilde{\alpha}}$. For $\frac{\bar{d}^{2}+\bar{d}}{2}$, we have $\bar{d}^{2}+\bar{d}\leq T^{2\tilde{\alpha}}+T^{\tilde{\alpha}}$ and $2\tilde{\alpha}\leq 1+\tilde{\alpha}$. Thus, we obtain $D=T^{1+\tilde{\alpha}}$. The remaining argument follows from the proof of Proposition 1. $\blacksquare$
Remark 2.
Theorem 1 indicates that when the total sampling number is small, i.e., when $a$ is small, the regret primarily stems from estimation error. In contrast, as the number of samples increases, the regret due to delay becomes more significant. Furthermore, Proposition 1 and Corollary 1 show that as the total number of samples over the horizon $T$ increases, the system exhibits a greater tolerance for delays.
@@ 3.2 Two-point risk-averse learning
When the function values can be queried at two action points at the same time slot, the perturbed actions are modified as $\hat{x}_{t}^{1}=x_{t}+\delta u_{t}$ and $\hat{x}_{t}^{2}=x_{t}-\delta u_{t}$. We query the stochastic cost for $m_{t}$ times under the actions $\hat{x}_{t}^{1}$ and $\hat{x}_{t}^{2}$, respectively, i.e., $J(\hat{x}_{t}^{1},\xi_{t}^{i})$ and $J(\hat{x}_{t}^{2},\xi_{t}^{j})$ for $i=1,\dots,m_{t}$ and $j=m_{t}+1,\dots,2m_{t}$. We adopt the same sampling strategy $m_{t}=\phi(t)$ as in (5). The performance of the algorithm is evaluated by the cumulative loss under the performed two actions against the best action in hindsight:
$\mathrm{R}_{2}(T)=\sum_{t=1}^{T}\frac{1}{2}\big{(}C_{t}(\hat{x}_{t}^{1})+C_{t}(\hat{x}_{t}^{2})\big{)}-\sum_{t=1}^{T}C_{t}(x^{\ast}).$ (23)
Assume that the samples generated simultaneously experience the same delay. At time slot $t$, the learner receives samples
$\mathcal{L}_{t}$ =$=\{J_{s_{n}|t}(\hat{x}_{s_{n}|t}^{1},\xi_{s_{n}|t}^{i}),J_{s_{n}|t}(\hat{x}_{s_{n}|t}^{2},\xi_{s_{n}|t}^{j}),u_{s_{n}|t}\ |$
@@
The empirical distribution functions constructed by the collected delayed feedback are given as
$\hat{F}_{s_{n}|t}^{1}(y)$ $=\frac{1}{m_{s_{n}}}\sum_{i=1}^{m_{s_{n}}}\mathbf{1}\{J_{s_{n}|t}(\hat{x}_{s_{n}|t}^{1},\xi_{s_{n}|t}^{i})\leq y\},$
$\hat{F}_{s_{n}|t}^{2}(y)$ $=\frac{1}{m_{s_{n}}}\sum_{i=m_{s_{n}}+1}^{2m_{s_{n}}}\mathbf{1}\{J_{s_{n}|t}(\hat{x}_{s_{n}|t}^{2},\xi_{s_{n}|t}^{i})\leq y\},$ (25)
@@
When the learner does not receive any loss at time slot $t$, we have $x_{t+1}=x_{t}$. Otherwise, the learning algorithm processes as
$\begin{array}[]{rl}x_{t}^{n}&=\mathcal{P}_{\mathcal{X}^{\delta}}(x_{t}^{n-1}-\eta\bar{g}_{t}^{n}),\ {\rm for}~{}n=1,\dots,\Delta_{t},\\ x_{t+1}&=x_{t}^{\Delta_{t}}.\end{array}$ (27)
The two-point risk-averse learning algorithm with delayed feedback is summarized in Algorithm 2.
@@
8:     collect feedback $\mathcal{L}_{t}$ as in (3.2)
@@
14:           build empirical distribution functions $\hat{F}_{s_{n}|t}^{1}(y)$ and $\hat{F}_{s_{n}|t}^{2}(y)$ as in (3.2)
@@
16:           estimate CVaR gradient $\bar{g}_{t}^{n}$ as in (26)
@@
The following theorem analyzes the regret bound achieved by Algorithm 2.
Theorem 2.
Let Assumptions 1, 2 and 3 hold. Let the sampling numbers over the iteration horizon $T$ satisfy (5). Selecting $\delta=T^{-\frac{a}{4}}$ and $\eta=(T+D)^{-\frac{1}{2}}$, Algorithm 1 achieves ${\rm R}_{2}(T)=\tilde{\mathcal{O}}\Big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{1}{2}}\Big{)}$ with high probability.
@@
${\rm R}_{2}(T)$ $=\sum_{t=1}^{T}\big{(}C_{t}(x_{t}+\delta u_{t})+C_{t}(x_{t}-\delta u_{t})\big{)}/2-\sum_{t=1}^{T}C_{t}(x^{\ast})$
$=\sum_{t=1}^{T}C_{t}(x_{t}+\delta u_{t})-\sum_{t=1}^{T}C_{t}(x^{\ast}),$ (28)
where the first equality is from the definition of $\hat{x}^{1}$ and $\hat{x}^{2}$. The second equality is from the symmetric distribution of $u_{t}$. The proof then follows from Theorem 1. The derivation differs starting from (3.1). The CVaR gradient estimate is bounded by
$\mathbb{E}[\|\bar{g}_{t}\|]$ $\leq\mathbb{E}_{u}\Big{[}\frac{d}{2\delta}\Big{\|}\big{(}C_{t}(x_{t}+\delta u_{t})-C_{t}(x_{t}-\delta u_{t})\big{)}\Big{\|}\big{\|}u_{t}\big{\|}\Big{]}$
@@
for all $t$, where the second inequality is from the Lipshcitzness of $C_{t}$. Similar to (3.1), we have $\mathbb{E}[\|\tilde{x}_{\tau-\tilde{s}_{\tau}}-\tilde{x}_{\tau}\|]\leq\mathbb{E}[\eta\tilde{s}_{\tau}\|\tilde{\bar{g}}_{\tau}\|]\leq L_{0}d\eta\tilde{s}_{\tau}$ with $\tilde{\bar{g}}_{\tau}=\bar{g}_{\mu(\tau)}$. Additionally, in the two-point case, constructing the empirical distribution function using finite samples induces CVaR gradient estimate error:
@@
$\hskip 10.00002pt-\big{(}{\rm CVaR}_{\alpha}[F_{t}^{1}]-{\rm CVaR}_{\alpha}[F_{t}^{2}]\big{)}.$ (30)
Denote $\tilde{\bar{\varepsilon}}_{\tau}=\bar{\varepsilon}_{\mu(\tau)}$, we have
$\hskip 10.00002pt\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(\tilde{x}_{\tau-\tilde{s}_{\tau}})-\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(\tilde{x}_{\tau})$
$\leq\sum_{\tau=1}^{L_{T}}\big{\langle}\nabla\tilde{C}_{\tau},\tilde{x}_{\tau-\tilde{s}_{\tau}}-\tilde{x}_{\tau}\big{\rangle}$
$=\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\big{\langle}\tilde{\bar{g}}_{\tau}-\frac{d}{\delta}\tilde{\bar{\varepsilon}}_{\tau}\tilde{u}_{\tau},\tilde{x}_{\tau-\tilde{s}_{\tau}}-\tilde{x}_{\tau}\big{\rangle}\Big{]}$
$\leq\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\|\tilde{\bar{g}}_{\tau}\|\|\tilde{x}_{\tau-\tilde{s}_{\tau}}-\tilde{x}_{\tau}\|+\|\frac{d}{\delta}{\tilde{\bar{\varepsilon}}_{\tau}}\tilde{u}_{\tau}\|\|\tilde{x}_{\tau-\tilde{s}_{\tau}}-\tilde{x}_{\tau}\|\Big{]}$
$\leq\sum_{\tau=1}^{L_{T}}\eta L_{0}^{2}d^{2}\tilde{s}_{\tau}+R_{21}\leq\eta L_{0}^{2}d^{2}D+R_{21},$ (31)
with $R_{21}=\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\|\frac{d}{\delta}{\tilde{\bar{\varepsilon}}_{\tau}}\tilde{u}_{\tau}\|\|\tilde{x}_{\tau-\tilde{s}_{\tau}}-\tilde{x}_{\tau}\|\Big{]}$. The first inequality is from the convexity of $C_{t}$ and the first equality is from substituting (3.2) into (3.2). The third inequality is from substituting (3.2) into (3.2). By the DKW inequality, we have that
@@
with probability at least $1-\mathbb{P}\{\bigcup_{i=1}^{2}B_{t}^{i}\}=1-2\bar{\gamma}$, for $t=1,\dots,L_{T}$. Let $\gamma=\bar{\gamma}L_{T}$. Then, by substituting (33) into $R_{21}$, we obtain
$R_{21}$ $\leq\frac{dD_{x}}{\delta}\sum_{\tau=1}^{L_{T}}\mathbb{E}[\|\tilde{\bar{\varepsilon}}_{\tau}\|]$
$\leq\frac{2dUD_{x}}{\alpha\delta}\sum_{\tau=1}^{L_{T}}\sqrt{\frac{\ln(2L_{T}/\gamma)}{2m_{t}}}$
$\leq\frac{2cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2L_{T}/\gamma)}{2}}L_{T}^{1-\frac{a}{2}},$ (34)
@@
with probability at least $1-2\gamma$, which establishes as $1-\mathbb{P}\{\bigcup_{\tau=1}^{T}\bigcup_{i=1}^{2}B_{\tau}^{i}\}\geq 1-\sum_{\tau=1}^{T}\sum_{i=1}^{2}\mathbb{P}\{B_{\tau}^{i}\}\geq 1-2\gamma$. Similar to (3.1), we have
$\hskip 10.00002pt\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(\tilde{x}_{\tau})-\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(x_{\delta}^{\ast})$
@@
$\leq\frac{D_{x}^{2}}{\eta}+\frac{\eta}{2}L_{0}^{2}d^{2}L_{T}+\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\|\frac{d}{\delta}{\tilde{\bar{\varepsilon}}_{\tau}}\tilde{u}_{\tau}\|\|\tilde{x}_{\tau}-x_{\delta}^{\ast}\|\Big{]}$
@@
$\leq\frac{D_{x}^{2}}{\eta}+\frac{\eta}{2}L_{0}^{2}d^{2}L_{T}+\frac{2cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2L_{T}/\gamma)}{2}}L_{T}^{1-\frac{a}{2}},$ (35)
@@
where the first inequality follows from (3.1) and $\mathbb{E}[\|\tilde{\bar{g}}_{\tau}\|^{2}]\leq L_{0}^{2}d^{2}$, the second inequality follows the derivation of $R_{21}$. Substituting (3.1), (3.2), (3.2) and (3.2) into (3.1), with $L_{T}\leq T$, we have
${\rm R}_{2}(T)$ $\leq(3+D_{x}/r)\delta L_{0}T+2U\bar{d}+\eta L_{0}^{2}d^{2}D+\frac{D_{x}^{2}}{\eta}$
$\hskip 10.00002pt+\frac{\eta}{2}L_{0}^{2}d^{2}T+\frac{2cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}}.$ (36)

### v2

@@ 3.1 One-point risk-averse learning
with probability $1-\gamma$. When $a<1$, selecting $\delta=T^{-\frac{a}{4}}\sqrt{\bar{d}P_{T}}$ and $\eta=(T+D)^{-\frac{3a}{4}}\sqrt{\bar{d}P_{T}}$, Algorithm 1 achieves ${\rm DR}_{1}(T)=\tilde{\mathcal{O}}\Big{(}\big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{3a}{4}}\big{)}\sqrt{\bar{d}P_{T}}+(T+D)^{-\frac{3a}{4}}T^{\frac{a}{4}}D\Big{)}$ with probability $1-\gamma$. When $a\geq 1$, selecting $\delta=T^{-\frac{1}{4}}\sqrt{\bar{d}P_{T}}$ and $\eta=(T+D)^{-\frac{3}{4}}\sqrt{\bar{d}P_{T}}$, Algorithm 1 achieves ${\rm DR}_{1}(T)=\tilde{\mathcal{O}}\Big{(}(T+D)^{\frac{3}{4}}\sqrt{\bar{d}P_{T}}\Big{)}$ with probability $1-\gamma$. $\blacksquare$
Remark 2
In Theorem 1, it can be observed $T^{1-\frac{a}{4}}$, $(T+D)^{\frac{3a}{4}}$ and $(T+D)^{-\frac{3a}{4}}T^{\frac{a}{4}}D$ are all smaller than $(T+D)^{1-\frac{a}{4}}$ when $a<1$. Thus, we conclude that $\tilde{{\mathrm{R}}}_{1}(T)=\tilde{\mathcal{O}}\Big{(}(T+D)^{1-\frac{a}{4}}\sqrt{\bar{d}P_{T}}\Big{)}$ is a more concise yet conservative compared to ${\rm DR}_{1}(T)$ in Theorem 1 when $a<1$. Additionally, it can be observed that the regret order of Algorithm 1 decreases as the sampling number parameter $a$ (i.e., the total number of samples over horizon $T$) increases and increases as the delay or path length of comparator increases.
Remark 3
The lower regret bound of one-point bandit learning in the delay-free case is $\tilde{\mathcal{O}}\Big{(}T^{\frac{3}{4}}\sqrt{1+P_{T}}\Big{)}$ [27]. In this work, we address the risk-averse learning and obtain a regret bound of $\tilde{\mathcal{O}}\big{(}(T+D)^{1-\frac{1}{4}\min{\{1,a\}}}\big{)}\sqrt{\bar{d}P_{T}}$. The dependence on the sampling parameter $a$ is from using queried cost values to construct the empirical distribution functions for the stochastic costs. When the number of samples is sufficiently large, i.e., $a$ is close to $1$, the term $(T+D)^{1-\frac{a}{4}}$ becomes $(T+D)^{\frac{3}{4}}$. In the absence of delay, the obtained regret bound then matches the result in [27]. Additionally, similar to [27], the dynamic regret bound exhibits a square-root dependence on the path-length, which becomes non-sublinear when $P_{T}\geq\sqrt{T}$. Nevertheless, in many practical cases, the path-length is relatively small.
@@ 3.2 Two-point risk-averse learning
In this section, we investigate the risk-averse learning using a two-point zeroth-order optimization method, which improves both the accuracy of CVaR estimation and the learning performance. At each time slot $t$, the stochastic costs are queried under two perturbed actions, i.e., $\hat{x}_{t}^{1}=x_{t}+\delta u_{t}$ and $\hat{x}_{t}^{2}=x_{t}-\delta u_{t}$, for $m_{t}$ times, which yields stochastic costs $J(\hat{x}_{t}^{1},\xi_{t}^{i})$ and $J(\hat{x}_{t}^{2},\xi_{t}^{j})$ for $i=1,\dots,m_{t}$ and $j=m_{t}+1,\dots,2m_{t}$. The sampling strategy $m_{t}=\phi(t)$ is the same as (6). Then, the performance of the algorithm is evaluated by the cumulative loss under the performed two actions against the best actions in hindsight, which is given by
$\mathrm{DR}_{2}(T)=\sum_{t=1}^{T}\frac{1}{2}\big{(}C_{t}(\hat{x}_{t}^{1})+C_{t}(\hat{x}_{t}^{2})\big{)}-\sum_{t=1}^{T}C_{t}(x_{t}^{\ast}).$ (30)
Assume that the samples generated simultaneously experience the same delay. At time slot $t$, the learner receives samples:
$\mathcal{L}_{t}$ $=\{J_{s_{n}|t}(\hat{x}_{s_{n}|t}^{1},\xi_{s_{n}|t}^{i}),J_{s_{n}|t}(\hat{x}_{s_{n}|t}^{2},\xi_{s_{n}|t}^{j}),u_{s_{n}|t}\ |$
@@
The empirical distribution functions constructed by the collected delayed feedback are given by
$\hskip-10.00002pt\hat{F}_{s_{n}|t}^{1}(y)$ $=\frac{1}{m_{s_{n}}}\sum_{i=1}^{m_{s_{n}}}\mathbf{1}\{J_{s_{n}|t}(\hat{x}_{s_{n}|t}^{1},\xi_{s_{n}|t}^{i})\leq y\},$
$\hskip-10.00002pt\hat{F}_{s_{n}|t}^{2}(y)$ $=\frac{1}{m_{s_{n}}}\sum_{j=m_{s_{n}}+1}^{2m_{s_{n}}}\mathbf{1}\{J_{s_{n}|t}(\hat{x}_{s_{n}|t}^{2},\xi_{s_{n}|t}^{j})\leq y\},$ (32)
@@
The learning process is the same as (11), where the gradient estimate is replaced with $\bar{g}_{t}^{n}$. The two-point risk-averse learning algorithm with delayed feedback is summarized in Algorithm 2.
@@
8:  Collect feedback $\mathcal{L}_{t}$, as in (3.2)
@@
14:    Build empirical distribution functions $\hat{F}_{s_{n}|t}^{1}(y)$ and $\hat{F}_{s_{n}|t}^{2}(y)$, as in (3.2)
@@
16:    Estimate CVaR gradient $\bar{g}_{t}^{n}$, as in (33)
@@
The following theorem analyzes the regret bound of Algorithm 2.
Theorem 2
Let Assumptions 1, 2 and 3 hold. Let the sampling numbers over the iteration horizon $T$ satisfy (6). Selecting $\delta=T^{-\frac{a}{4}}\sqrt{\bar{d}P_{T}}$ and $\eta=(T+D)^{-\frac{1}{2}}\sqrt{\bar{d}P_{T}}$, Algorithm 2 achieves ${\rm DR}_{2}(T)=\tilde{\mathcal{O}}\bigg{(}\big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{1}{2}}\big{)}\sqrt{\bar{d}P_{T}}\bigg{)}$ with high probability.
@@
${\rm DR}_{2}(T)$ $\hskip-1.99997pt=\hskip-1.99997pt\sum_{t=1}^{T}\big{(}C_{t}(x_{t}+\delta u_{t})+C_{t}(x_{t}-\delta u_{t})\big{)}/2-\hskip-3.00003pt\sum_{t=1}^{T}C_{t}(x_{t}^{\ast})$
$=\sum_{t=1}^{T}C_{t}(x_{t}+\delta u_{t})-\sum_{t=1}^{T}C_{t}(x_{t}^{\ast}),$ (34)
where the first equality is from the definition of $\hat{x}^{1}$ and $\hat{x}^{2}$. The second equality is from the symmetric distribution of $u_{t}$. The proof follows the same approach as in Theorem 1. The derivation differs starting from (3.1). The CVaR gradient estimate (33) is bounded by
$\mathbb{E}[\|\bar{g}_{t}\|]$ $\leq\mathbb{E}_{u}\Big{[}\frac{d}{2\delta}\Big{\|}C_{t}(x_{t}+\delta u_{t})-C_{t}(x_{t}-\delta u_{t})\Big{\|}\big{\|}u_{t}\big{\|}\Big{]}$
@@
for all $t$, where the second inequality is from the Lipschitzness of $C_{t}$. Similar to (3.1), we have
$\|x_{\mu(\tau)}-\tilde{x}_{\tau}\|\leq\sum_{j=0}^{\tilde{s}_{\tau}-1}\eta\|\tilde{\bar{g}}_{\tau-\tilde{s}_{\tau}+j}\|\leq L_{0}d\eta\tilde{s}_{\tau}$ (36)
where $\tilde{\bar{g}}_{i}$ is $i$-th utilized two-point CVaR gradient estimate. Thus, for $R_{1}$ in (3.1), we have
$R_{1}$ $\leq\sum_{\tau=1}^{L_{T}}L_{0}\|x_{\mu(\tau)}-\tilde{x}_{\tau}\|\leq L_{0}^{2}d\eta\sum_{\tau=1}^{L_{T}}\tilde{s}_{\tau}=L_{0}^{2}d\eta D,$ (37)
where the first inequality is from the Lipschitzness of $C_{t}$. The second inequality is from substituting (36) into (37). The last inequality is from the definition of cumulative delay as in (12).
Additionally, in the two-point case, the error of the CVaR gradient estimate (33) is given as
@@
$\hskip 10.00002pt-\big{(}{\rm CVaR}_{\alpha}[F_{t}^{1}]-{\rm CVaR}_{\alpha}[F_{t}^{2}]\big{)},$ (38)
where $\hat{F}_{t}^{1}$ and $\hat{F}_{t}^{2}$ the empirical distribution functions at time $t$ and $F_{t}^{1}$ and $F_{t}^{2}$ denote the corresponding true distribution function. By the DKW inequality, we have that
@@
with probability at least $1-\mathbb{P}\{\bigcup_{i=1}^{2}B_{t}^{i}\}=1-2\bar{\gamma}$, for $t=1,\dots,L_{T}$. Similar to (3.1), we have
$R_{2}$ $\leq R_{21}+\bar{R}_{22}+\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\frac{\eta}{2}\|\tilde{\bar{g}}_{\tau}\|^{2}\Big{]}$
$\leq R_{21}+\bar{R}_{22}+\frac{\eta}{2}L_{0}^{2}d^{2}L_{T}.$ (41)
@@
with $\bar{R}_{22}=\sum_{\tau=1}^{L_{T}}\mathbb{E}[\|\frac{d}{\delta}{\tilde{\bar{\varepsilon}}_{\tau}}\tilde{u}_{\tau}\|\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|]$ and $\tilde{\bar{\varepsilon}}_{\tau}=\bar{\varepsilon}_{\mu(\tau)}$. Let $\gamma=\bar{\gamma}L_{T}$. Then, we obtain
@@
$\bar{R}_{22}$ $\leq\frac{dD_{x}}{\delta}\sum_{\tau=1}^{L_{T}}\mathbb{E}[|\tilde{\bar{\varepsilon}}_{\tau}|]\leq\frac{2dUD_{x}}{\alpha\delta}\sum_{\tau=1}^{L_{T}}\sqrt{\frac{\ln(2L_{T}/\gamma)}{2\tilde{m}_{\tau}}}$
$\leq\frac{2cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}},$ (42)
with probability at least $1-2\gamma$, which establishes as $1-\mathbb{P}\{\bigcup_{\tau=1}^{T}\bigcup_{i=1}^{2}B_{\tau}^{i}\}\geq 1-\sum_{\tau=1}^{T}\sum_{i=1}^{2}\mathbb{P}\{B_{\tau}^{i}\}\geq 1-2\gamma$. The second inequality is from substituting (40) into (3.2).
Similar to (3.1)-(28), we have
$R_{3}$ $\leq\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}(\|\tilde{\bar{g}}_{\tau}\|+\|\frac{d}{\delta}\tilde{\bar{\varepsilon}}_{\tau}\tilde{u}_{\tau}\|)\|\tilde{x}_{\tau}^{\delta,\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\delta,\ast}\|\Big{]}$
$\leq 2\frac{dU}{\delta}\bar{d}P_{T}+\frac{2cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}}$ (43)
@@
where the second inequality follows the derivation of (27) and (28). Substituting (3.1), (3.1), (37), (3.2), (3.2) and (3.2) into (3.2), and with $L_{T}\leq T$, we have
${\rm DR}_{2}(T)$ $\leq(3+D_{x}/r)\delta L_{0}T+2U\bar{d}+L_{0}^{2}d\eta D+\frac{\eta}{2}L_{0}^{2}d^{2}T$
$\hskip-35.00005pt+\hskip-3.00003pt\frac{3D_{x}^{2}+4D_{x}\bar{d}P_{T}}{2\eta}\hskip-1.99997pt+\hskip-1.99997pt\frac{2dU\bar{d}P_{T}}{\delta}\hskip-1.99997pt+\hskip-1.99997pt\frac{4cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}}.$ (44)